OpenStudy (anonymous):

What is the equation of a circle with diameter AB that has endpoints A(0, 0) and B(8, 6)? WHOEVER ANSWERS GETS A MEDAL

OpenStudy (agent0smith):

First you need: the distance between (0, 0) and (8, 6). Use the distance formula. The midpoint between (0, 0) and (8, 6). Use the midpoint formula.

OpenStudy (anonymous):

is the distance 10?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

and the midpoint 4,3

OpenStudy (agent0smith):

(x - a)^2 + (y - b)^2 = r^2 Your centre (a, b) is (4, 3). Your radius r, is half of the diameter 10.

OpenStudy (anonymous):

so would it be (x - 3)2 + (y - 4)2 = 5 or (x - 4)2 + (y - 3)2 = 5

OpenStudy (agent0smith):

look at the a and b: your centre (a, b) is (4, 3) And it's r^2, not just r.